(8x-x^2y)dy+(x-xy^2)dx=0solve
1
Expert's answer
2021-08-04T18:13:10-0400

Given that

"(8x-x^2y)dy+(x-xy^2)dx=0"

There is no specified method mentioned in the question. So, I used exact method to solve this.

Taking "x-xy" as common

"(x-xy)((8-x)dy+(1-y)dx)=0"

"(8-x)dy+(1-y)dx=0"

"(1-y)dx+(8-x)dy=0"

The above equation is in the form of

"M(x,y)dx+N(x,y)dy=0"

"M(x,y)=(1-y)" and "N(x,y)=(8-x)"

"\frac{\partial M}{\partial y}=-1" and "\frac{\partial N}{\partial x}=-1"

"\frac{\partial M}{\partial y}=\frac{\partial N}{\partial x}"

So, the equation is exact.

So, the general solution is

"\int Mdx+\int (" terms of N not involving x)"dy=c" (here c is the constant)

"\int (1-y)dx +\int 8dy=c"

"x-xy+8y=c"

So, the solution is "x-xy+8y=c"